OpenStudy (bigbosssaint21):

Solve x^2 - 8x - 20 = 0. x = 4, x = -5 x = -4, x = 5 x = 2, x = -10 x = -2, x = 10

OpenStudy (imstuck):

Do you know how to factor this?

OpenStudy (imstuck):

Factoring it will give you the zeros of the polynomial (which is what you get with your answer choices from above)

OpenStudy (bigbosssaint21):

i dont know about factoring =/

OpenStudy (mathmale):

Factoring should work well. Alternative approaches include the quadratic formula and graphing. You have a quadratic equation: x^2 - 8x - 20 = 0. Look at the constant term, -20. What are some of the possible pairs of factors of that? for example, -20 = (-4)(5). But that pair must add up to the coefficient of x in the quadratic equation. Does it? No. So you'll have to find another set of factors of -20 and determine whether the sum of those two factors adds up to equal the coefficient of x. Try it, please.

OpenStudy (mathmale):

+20 factors into the following pairs: (1)(20), (2)(10), (4)(5). 1+20 =21 2+10=12 4+5 = 9 Would you please do the same thing for -20? -20 factors into (-1)(20), (and what other pairs?)